Corner Coffee Shop’s latte’s are $3. You are a barista at a different coffee shop, and you know that it takes about $0.5 worth of inputs to make a latte.

A.) What is Corner Coffee Shop’s % markup?

B.) What is Corner Coffee Shop’s % margin?
Answer #1

The price of a latte is $3 and the cost is $0.5. Now markup is defined as (sale price-cost of sales)/cost of sales

Thus, the coffee shops' markup =(3-0.5)/0.5*100=500%.

The margin is defined as (sale price-cost of sales)/price of sales

The coffee shop's margin = (3-0.5)/3*100=83.33%